80 liked lemon lime.
68 liked cherry cola.
42 liked both types.
a) How many liked only lemon lime?
b) How many liked only cherry cola?
c) How many liked either one or the other or both?
d) How many did not like either flavor?
out of 125 people
need answer for c and d only thank you

Found 2 solutions by edjones, jorel1380:
Answer by edjones(8007)

Make a Venn diagram to understand this.
a)80-42=38
b)68-42=26
c)38+42+26=106 (A U B)
d)125-106=19 (A U B)'
.
Ed

Answer by jorel1380(3719)

80+68=148
c.148-42=106 liked either lemon lime or cherry cola or both.
d.125-106=19 liked neither lemon lime or cherry cola
a.80-42=38 liked lemon lime only
b.68-42=26 liked cherry only